2.) How long would it take for a current of 100A to deposit 10 grams of Fe0 from a solution of FeCl2.
1
Expert's answer
2020-10-05T14:27:55-0400

From Faraday's law:

m=ItF×M/z, where F is a Faraday's constant, z is a valency, M is a molar mass.

t=mFz/MI,

t=10×96485×2/72×100=268(sec).